ROTATIONAL SPECTROSCOPY OF THE NH_{3}-SO_{3}-H_{2}O SYSTEM

\underline{S. W. HUNT}, C. S. BRAUER, M. CRADDOCK, K. R. LEOPOLD, Department of Chemistry, University of Minnesota, Minneapolis, MN 55455.

Previous reports of atmospheric sulfate aerosol formation indicate significantly enhanced nucleation rates in the presence of elevated concentrations of ammonia.^{a} Computational studies^{b} predict the most stable structure for NH_{3}-SO_{3}- H_{2}O ternary system is the hydrogen bonded complex, H_{2}SO_{4}-NH_{3}, but a second minimum, only 0.4 kcal/mol higher in energy occurs for the hydrated form of sulfamic acid (H_{3}N-SO_{3})·H_{2}O. In this report, we present microwave spectroscopic results obtained in a supersonic jet seeded with H_{2}O, NH_{3}, and SO_{3}. A-type rotational spectra of a near-prolate asymmetric top requiring the simultaneous presence of NH_{3}, H_{2}O, and SO_{3} in the jet have been observed. The transitions correlate with the intensity of known H_{3}N-SO_{3} spectra, and spectroscopic constants are consistent with those predicted for (H_{3}N-SO_{3})·H_{2}O.^{b} Assignment to H_{2}SO_{4}-NH_{3}, however, cannot be ruled out on the basis of data corresponding to the parent isotopic form alone. Isotopic work is in progress, and spectral results will be reported. A novel, dual-injection arrangement, in which two gases are independently admitted into the early phases of the supersonic expansion, has been developed and will be described as well.

aR. J. Werber; P. H. McMurry; L. Mauldin; D. J. Tanner; F. L. Eisle; F. J. Brechtel; S. M. Kreidenweis; G. L. Kok; R. D. Schillawski; D. Baumgardner J. Geophys Res. 103, 16385 (1998).

bL. J. Larson; F-M. Tao, J. Phys. Chem. A, in press.